i need help on solving for elmination, i get confused.
Problem: 2x+6y=17
2x-10y=9
break it down for me step by step, if you can..thank you! :)

Answer by edjones(8007)

2x+6y=17
2x-10y=9
----------------subtract
==>16y=8
y=1/2
.
2x+3=17 replace y with 1/2
2x=14
x=7
.
Ed